OpenStudy (anonymous):

How do you find the unit cost for a) 4l of milk for 4.29

hartnn (hartnn):

unit cost means cost of 1 litre of milk. you have 4l milk costing 4.29 so, to get the cost of 1l of milk, just divide the amount by 4, what you get ?

OpenStudy (unklerhaukus):

\[\frac{4.29}{4}\frac{[$]}{[\text L]}=\left(\frac{4}{4}+\frac{0.29}4\right)\frac{[$]}{[\text L]}\]\[\qquad=\left(1+\frac{0.24}4+\frac{0.05}{4}\right)\frac{[$]}{[\text L]}\]\[\quad=\left(1+0.06+\frac{0.04}{4}+\frac{0.01}4\right)\frac{[$]}{[\text L]}\]\[=\Big(1.06+0.01+0.0025\Big)\frac{[$]}{[\text L]}\]\[=\Big(1.06+0.0125\Big)\frac{[$]}{[\text L]}\]\[=\]